OpenStudy (anonymous):

Consider the functions f(x) = 13x2 and g(x) = 7x − 5. What is the value of (f + g)(5)?

hero (hero):

Hint: (f+g)(x) = f(x) + g(x)

OpenStudy (anonymous):

thanks i got it

hero (hero):

What did you get for the final result?

hero (hero):

You calculated f(5) + g(5) ?

OpenStudy (anonymous):

355

hero (hero):

13(5)^2 + 7(5) - 5 = 355 True

OpenStudy (anonymous):

Thank you for the help